OpenStudy (anonymous):

A pillar candle is shaped like a cylinder with a radius of 1.5 inches. The candle holds 123 cubic inches of wax. How tall is the candle? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (studygurl14):

Formula for volume of a cylinder: \(\large V=r^2h\pi\) Plug in: \(\large 123=(1.5)^2h(3.14)\) Solve for h

OpenStudy (xapproachesinfinity):

it is really just about using the volume formula for a cylinder and solve for h